For the cashflows, the function for evaluating the ROR is:

(Cashflow)
Year Cashflow
0 -$500
1 0
2 0
3 400
4 -50
5 300

a. -500 - 50(P/F, i, 4) + 400(P/F, i, 3) + 300 (P/F,i,5) = 0
b. -500 - 50(P/F, i, 2) + 400(P/F, i, 1) + 300 (P/F,i,3) = 0
c. 500 - 50(P/F, i, 4) + 400(P/F, i, 3) + 300 (P/F,i,5) = 0
d. None of the above
1
Expert's answer
2015-11-04T00:00:47-0500
Year Cashflow
0 -$500
1 0
2 0
3 400
4 -50
5 300
For the cashflows, the function for evaluating the ROR is:
ROR = -500 - 50(P/F, i, 4) + 400(P/F, i, 3) + 300 (P/F,i,5) = 0

So, the right answer is a. -500 - 50(P/F, i, 4) + 400(P/F, i, 3) + 300 (P/F,i,5) = 0